\begin{thebibliography}{99}
\bibitem{aim} AIM Minimum Rank - Special Graphs Work Group (F. Barioli, W. Barrett, S. Butler, S. Cioaba, D. Cvetkovic, S. Fallat, C. Godsil, W. Haemers, L. Hogben, R. Mikkelson, S. Narayan, O. Pryporova, I. Sciriha, W. So, D. Stevanovic, H. van der Holst, K.V. Meulen, A.W. Wehe), Zero forcing sets and the minimum rank of graphs, Linear Algebra and its Applications 428 (2008) 1628-1648.
\bibitem{ahl} N. Alon, S. Hoory, N. Linial, The Moore Bound for Irregular Graphs, Graphs and Combinatorics 18 (2002) 53-57.
\bibitem{aim2} F. Barioli, W. Barrett, S. Fallat, H.T. Hall, L. Hogben, B. Shader, P. van den Driessche, H. van der Holst, Parameters related to tree-width, zero forcing, and maximum nullity of a graph, Journal of Graph Theory 72 (2013) 146-177.
\bibitem{bg} D. Burgarth, V. Giovannetti, Full control by locally induced relaxation, Physical Review Letters 99 (2007) 100501.
\bibitem{bghsy} D. Burgarth, V. Giovannetti, L. Hogben, S. Severini, M. Young, Logic circuits from zero forcing, Natural Computing 14 (2015) 485-490.
\bibitem{bm} D. Burgarth, K. Maruyama, Indirect Hamiltonian identification through a small gateway, New Journal of Physics 11 (2009) 103019.
\bibitem{cs} L.S. Chandran, C.R. Subramanian, Girth and treewidth, Journal of Combinatorial Theory, Series B 93 (2005) 23-32.
\bibitem{dh} R. Davila, M. Henning, The Forcing Number of Graphs with Given Girth, arXiv:1610.08435v1.
\bibitem{dks} R. Davila, T. Kalinowski, S. Stephen, Proof of a conjecture of Davila and Kenter regarding a lower bound for the forcing number in terms of girth and minimum degree, arXiv:1611.06557v2.
\bibitem{dk} R. Davila, F. Kenter, Bounds for the Zero-Forcing Number of Graphs with Large Girth, arXiv:1406.0482v2.
\bibitem{gprs} M. Gentner, L.D. Penso, D. Rautenbach, U.S. Souza, Extremal Values and Bounds for the Zero Forcing Number, Discrete Applied Mathematics 214 (2016) 196-200.
\bibitem{gr} M. Gentner, D. Rautenbach, Some Bounds on the Zero Forcing Number of a Graph, arXiv:1608.00747.
\bibitem{hv} L. Hogben, H. van der Holst, Forbidden minors for the class of graphs $G$ with $\zeta(G)\leq 2$, Linear Algebra and Its Applications 423 (2007) 42-52.
\bibitem{s} S. Severini, Nondiscriminatory propogation on trees, Journal of Physics A: Mathematical and Theoretical 41 (2008) 482002.
%\bibitem{td} M. Trefois, J.-C. Delvenne, Zero forcing number, constraint matchings and strong structural controllability, arXiv:1405.6222v1.
\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{latexsym,amssymb,amsmath,geometry,cite,enumerate,float,verbatim}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\usepackage{lineno}
\usepackage{setspace}
\allowdisplaybreaks

\begin{document}

\onehalfspace

\title{A Short Proof for a Lower Bound on the Zero Forcing Number}

\author{M. F\"{u}rst\and D. Rautenbach}

\date{}

\maketitle

\begin{center}
{\small 
Institute of Optimization and Operations Research, Ulm University, Germany\\
\texttt{maximilian.fuerst,dieter.rautenbach@uni-ulm.de}}
\end{center}

\begin{abstract}
We provide a short proof of a conjecture of Davila and Kenter concerning a lower bound on the zero forcing number $Z(G)$ of a graph $G$. 
More specifically, we show that $Z(G)\geq (g-2)(\delta-2)+2$
for every graph $G$ of girth $g$ at least $3$ and minimum degree $\delta$ at least $2$.
\end{abstract}

{\small 

\begin{tabular}{lp{13cm}}
{\bf Keywords:} Zero forcing; girth; Moore bound
\end{tabular}
}

\section{Introduction}

We consider finite, simple, and undirected graphs and use standard terminology.

For an integer $n$, let $[n]$ denote the set of positive integers at most $n$.
For a graph $G$,
a set $Z$ of vertices of $G$ is a {\it zero forcing set} of $G$ 
if the elements of $V(G)\setminus Z$ have a linear order $u_1,\ldots,u_k$ 
such that, for every $i$ in $[k]$, 
there is some vertex $v_i$ in $Z\cup \{ u_j:j\in [i-1]\}$ such that 
$u_i$ is the only neighbor of $v_i$ outside of $Z\cup \{ u_j:j\in [i-1]\}$;
in particular, 
$N_G[v_i]\setminus (Z\cup N_G[v_1]\cup \cdots \cup N_G[v_{i-1}])=\{ u_i\}$ for $i\in [k]$.
The {\it zero forcing number $Z(G)$} of $G$, 
defined as the minimum order of a zero forcing set of $G$,
was proposed by the AIM Minimum Rank - Special Graphs Work Group \cite{aim,hv} 
as an upper bound on the corank of matrices associated with a given graph.
The same parameter was also considered in connection with 
quantum physics \cite{bg,bm,s} and logic circuits \cite{bghsy}.

In \cite{dk} Davila and Kenter conjectured that 
\begin{eqnarray}
Z(G)&\geq &(g-2)(\delta-2)+2\label{eb3}
\end{eqnarray}
for every graph $G$ of girth $g$ at least $3$ and minimum degree $\delta$ at least $2$.
They observe that, for $g>6$ and sufficiently large $\delta$ in terms of $g$,
the conjectured bound follows by combining results from \cite{aim2} and \cite{cs}.
For $g\leq 6$, it was shown in \cite{gprs,gr},
Davila and Henning \cite{dh} showed it for $7\leq g\leq 10$,
and, eventually, 
Davila, Kalinowski, and Stephen \cite{dks} completed the proof.
The proof in \cite{dks} is rather short itself but relies on \cite{gprs,gr,dh}.
While the cases $g\leq 6$ have rather short proofs,
the proof in \cite{dh} for $7\leq g\leq 10$ extends over more than eleven pages and requires a detailed case analysis.
Therefore, the complete proof of (\ref{eb3}) obtained by combining \cite{gprs,gr,dh,dks} is rather long.

In the present note we propose a considerably shorter and simpler proof.
Our approach only requires a special treatment for the triangle-free case $g=4$ \cite{gprs},
involves a new lower bound on the zero forcing number,
and an application of the Moore bound \cite{ahl}.

\section{Proof of (\ref{eb3})}

Our first result is a natural generalization of the well known fact $Z(G)\geq \delta(G)$ \cite{aim},
where $\delta(G)$ is the minimum degree of a graph $G$.
For a set $X$ of vertices of a graph $G$ of order $n$, let 
$N_G(X)=\left(\bigcup\limits_{u\in X}N_G(u)\right)\setminus X$,
$N_G[X]=X\cup N_G(X)$, and
$\delta_p(G)=\min\left\{ |N_G(X)|:X\subseteq V(G)\mbox{ and }|X|=p\right\}$
for $p\in [n]$.
Note that $\delta_1(G)$ equals $\delta(G)$.

\begin{lemma}\label{lemma1}
If $G$ is a graph of order $n$, then $Z(G)\geq \delta_p(G)$ for every $p\in [n]$.
\end{lemma}
{\it Proof:} Let $Z$ be a zero forcing set of minimum order.
Let $u_1,\ldots,u_k$ and $v_1,\ldots,v_k$ be as in the introduction.
Since, by definition, $\delta_p(G)\leq n-p$, the result is trivial for $p\geq k=n-|Z|$,
and we may assume that $p<k$.
As noted above, we have $N_G[v_i]\setminus (Z\cup N_G[v_1]\cup \cdots \cup N_G[v_{i-1}])=\{ u_i\}$ for $i\in [k]$,
which implies that $X=\{ v_1,\ldots,v_p\}$ is a set of $p$ distinct vertices of $G$.
Furthermore, it implies that $|N_G[X]|\leq |Z|+p$,
and, hence, $\delta_p(G)\leq |N_G(X)|=|N_G[X]|-p\leq |Z|$ as required. $\Box$

\medskip

\noindent For later reference, we recall the Moore bound for irregular graphs.

\begin{theorem}[Alon, Hoory, and Linial \cite{ahl}]\label{theoremmoore}
If $G$ is a graph of order $n$, girth at least $2r$ for some integer $r$, and average degree $d$ at least $2$, then
$n\geq 2\sum\limits_{i=0}^{r-1}(d-1)^i$.
\end{theorem}
We also need the following numerical fact.

\begin{lemma}\label{lemma2}
For positive integers $p$ and $f$ with $p\geq 5$ and $2p-1\leq f\leq {p\choose 2}$, 
$$\left(1+\frac{2(f-p)}{f+p}\right)^{\lceil\frac{p}{2}\rceil+1}>f-p+1.$$
\end{lemma}
{\it Proof:} For $p\geq 17$,
it follows from $f\geq 2p-1$
that $1+\frac{2(f-p)}{f+p}\geq 1.64$, 
and, since $1.64^{\lceil\frac{p}{2}\rceil+1}>{p\choose 2}-p+1$,
the desired inequality follows for these values of $p$.
For the finitely many pairs $(p,f)$ with $5\leq p \leq 16$ and $2p-1\leq f\leq {p\choose 2}$, we verified it using a computer. 
$\Box$

\medskip

\noindent We proceed to the proof of (\ref{eb3}).

\begin{theorem}
If $G$ is a graph of girth $g$ at least $3$ and minimum degree $\delta$ at least $2$,
then $Z(G)\geq (g-2)(\delta-2)+2$.
\end{theorem}
{\it Proof:} For $g=3$, the inequality simplifies to the known fact $Z(G)\geq \delta(G)$,
and, for $g=4$, it has been shown in \cite{gprs}. Now, let $g\geq 5$.
Let $X$ be a set of $g-2$ vertices of $G$ with $|N_G(X)|=\delta_{g-2}(G)$,
and, let $N=N_G(X)$.
By the girth condition, the components of $G[X]$ are trees,
and no vertex in $N$ has more than one neighbor in any component of $G[X]$.

Let $K_1,\ldots,K_p$ be the vertex sets of the components of $G[X]$.

If $p\geq 3$, and there are two vertices in $N$
that both have neighbors in two distinct components of $G[X]$,
then $G$ contains a cycle of order at most $2+|K_i|+|K_j|\leq 2+(g-2)-(m-2)<g$
which is a contradiction.
Similarly, if $p=2$, and there are three vertices $u$, $v$, and $w$ in $N$
that both have neighbors in $K_1$ and $K_2$,
then let $u_i$, $v_i$, and $w_i$ denote the corresponding neighbors in $K_i$ for $i\in [2]$, respectively.
If $G[K_1]$ contains a path between two of the vertices $u_1$, $v_1$, and $w_1$
avoiding the third, then $G$ contains a cycle of order at most $2+(|K_1|-1)+|K_2|=g-1$, which is a contradiction.
By symmetry, this implies $u_1=v_1=w_1$ and $u_2=v_2=w_2$,
and $G$ contains the cycle $u_1uu_2vu_1$ of order $4$, which is a contradiction.

Combining these observations, we obtain
\begin{eqnarray}\label{e1}
\sum\limits_{\{ i, j\}\in {[p]\choose 2}}|N_G(K_i)\cap N_G(K_j)|
\leq 
\begin{cases}
{p\choose 2} & \mbox{, for $p\geq 3$, and }\\
2p-2 & \mbox{, for $p\in \{ 1,2\}$}.
\end{cases}
\end{eqnarray}
Let the bipartite graph $H$ arise from $G[X\cup N]$ 
by contracting each component of $G[X]$ to a single vertex, 
and removing all edges of $G[N]$.
Lemma \ref{lemma1} and a simple counting implies
\begin{eqnarray*}
Z(G) & \geq & \delta_{g-2}(G)\\ 
& = & |N|\\
& = & \sum_{u\in V(H)\setminus N}d_H(u)-\sum_{v\in N}(d_H(v)-1)\\
& \geq & \sum_{i=1}^p\Big(\delta|K_i|-2(|K_i|-1)\Big)-\sum_{v\in N}(d_H(v)-1)\\
& = & (g-2)(\delta-2)+2+\left((2p-2)-\sum_{v\in N}(d_H(v)-1)\right).
\end{eqnarray*}
In view of (\ref{eb3}), we may assume $f\geq 2p-1$ for $f=\sum\limits_{v\in N}(d_H(v)-1)$.

Since 
$$2p-1\leq f=\sum\limits_{v\in N}(d_H(v)-1)
\leq \sum\limits_{v\in N}{d_H(v)\choose 2}
=\sum\limits_{\{ i, j\}\in {[p]\choose 2}}|N_G(K_i)\cap N_G(K_j)|,$$
(\ref{e1}) implies $p\geq 5$.

Let $H'$ arise by removing all vertices of degree $1$ from $H$.
Since every vertex $u$ in $V(H)\setminus N$ satisfies
$d_H(u)\geq \delta |K_i|-2(|K_i|-1)\geq 2$ for some $i\in [p]$,
the graph $H'$ contains all $p$ vertices of $V(H)\setminus N$.
Let $H'$ contain $q$ vertices of $N$.
Since $H'$ has order $p+q$ and size
$$\sum\limits_{v\in N\cap V(H')}d_H(v)
=q+\sum\limits_{v\in N}(d_H(v)-1)=q+f,$$
its average degree is at least $\frac{2(f+q)}{p+q}$,
which is at least $2$, because $f\geq 2p-1\geq p$.

If $H'$ contains a cycle of order $2\ell$,
then $G$ contains a cycle of order at most $(g-2)-(p-\ell)+\ell$.
By the girth condition, 
this implies that the bipartite graph $H'$ has girth at least 
$p+2$, if $p$ is even, and $p+3$, if $p$ is odd.

Using Theorem \ref{theoremmoore} and $f\geq q$, we obtain
\begin{eqnarray*}
p+q & \geq & 2\sum\limits_{i=0}^{\lceil\frac{p}{2}\rceil}\left(\frac{2(f+q)}{p+q}-1\right)^i\\
&=& 2\frac{p+q}{2f-2p}\left(\left(1+\frac{2(f-p)}{p+q}\right)^{\lceil\frac{p}{2}\rceil+1}-1\right)\\
&\geq &2\frac{p+q}{2f-2p}\left(\left(1+\frac{2(f-p)}{p+f}\right)^{\lceil\frac{p}{2}\rceil+1}-1\right),
\end{eqnarray*}
which implies
$\left(1+\frac{2(f-p)}{f+p}\right)^{\lceil\frac{p}{2}\rceil+1}\leq f-p+1$.
Since $f\geq 2p-1$, and, by (\ref{e1}), $f\leq {p\choose 2}$,
this contradicts Lemma \ref{lemma2},
which completes the proof.
$\Box$

\end{document}